OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2).

OpenStudy (anonymous):

find your slope first, then take one of the points e.x (4, 8) and then plug m, x, and y into y=mx+b to solve for b

OpenStudy (anonymous):

that didn't help

OpenStudy (anonymous):

can you solve for the slope on your own first?

OpenStudy (anonymous):

no, i have no idea how to do slopes

OpenStudy (anonymous):

y = 5/2 x - 2 y = -5/2 x + 8 Y = 2/5 x - 6 y = -2/5 x + 4 these are the answers i get to choose from

OpenStudy (anonymous):